# Large Language Model Inference Acceleration: A Comprehensive Hardware Perspective

### v3

@@ 3.1.3. GPU
Weight-Activation Quantization. In addition to hardware units that support FP16 computations, NVIDIA GPUs also provide hardware units that support INT4, INT8, and FP8 computations. The number of these computation units can be 2$\times$ and 4$\times$ greater than FP16 on each chip. Compared to weight-only quantization, weight-activation quantization can utilize INT4, INT8, and FP8 computations, thereby maximizing the peak computational performance of the GPU. Since the prefill phase in LLM inference is compute-bound, weight-activation quantization can significantly enhance performance during this stage. LLM.int8 (dettmers2022gpt3, ) uses vector-wise quantization with separate normalization constants for each inner product in the matrix multiplication, to quantize most of the features. For the outliers, it isolates the outlier feature dimensions into a 16-bit matrix multiplication while still more than 99.9% of values are multiplied in 8-bit. For BLOOM-176B model, LLM.int8 achieves 4.05 tokens/s, 30.3 tokens/s and 109.77 tokens/s for batch size 1, 8 and 32, respectively, on 3 A100 GPUs in decode phase. The inference speed is slightly slower but close to 16-bit inference with less GPU consumption. SmoothQuant (xiao2023smoothquant, ) enables 8-bit weight and 8-bit activation (W8A8) quantization for LLMs. Based on the fact that weights are easy to quantize while activations are not, SmoothQuant smooths the activation outliers by offline migrating the quantization difficulty from activations to weights with a mathematically equivalent transformation. SmoothQuant enables an INT8 quantization of both weights and activations for all the matrix multiplications in LLMs. SmoothQuant achieves up to 1.56$\times$ speedup and 2$\times$ memory reduction for LLMs with negligible loss in accuracy. QUIK (ashkboos2023towards, ) is for the first time, that the majority of inference computations for LLMs can be performed with both weights and activations being cast to 4 bits. QUIK compresses most of the weights and activations to 4-bit, while keeping some outlier weights and activations in higher-precision. It also provides GPU kernels matching the QUIK format with highly-efficient layer-wise runtimes, which lead to practical end-to-end throughput improvements of up to 3.4$\times$ relative to FP16 execution in prefill phase.
@@
Prevalent quantization schemes (e.g., W8A8) cannot fully leverage the capabilities of modern GPUs, such as 4-bit integer operators, resulting in sub-optimal performance. To maximize the throughput, Atom (zhao2024atom, ) significantly boosts serving throughput by using low-bit operators and considerably reduces memory consumption via low-bit quantization. It attains high accuracy by applying a novel mixed-precision and fine-grained quantization process. For single batch inference, Atom can achieve about 30 tokens/s for on a NVIDIA RTX 4090 GPU. Atom improves end-to-end throughput by up to 7.73$\times$ compared to the FP16 and by 2.53$\times$ compared to INT8 quantization, while maintaining the same latency target.
@@
Compared to integer quantization, floating-point (FP) quantization can better handle long-tail or bell-shaped distributions, and it has emerged as a default choice in many hardware platforms. LLM-FP4 (liu2023llm, ) quantizes both weights and activations in LLMs down to 4-bit floating-point values (W4A4) with negligible accuracy loss. Due to the lack of PF4 computing unit in GPUs, its decoding speed maybe slower than FP16 baseline.
3.1.4. FPGA
@@
Weight-Activation Quantization. FlexRun (hur2023fast, ) uses 8-bit quantization (W8A8), conducts an in-depth design space exploration to find the best accelerator architecture for a target LLM model, and automatically reconfigures the accelerator based on the exploration results. With the implementation on Intel Stratix 10 GX and MX FPGAs, FlexRun outperforms the current state-of-the-art FPGA-based accelerator by 1.15$\times$–1.50$\times$ for GPT2, respectively. Compared to Nvidia’s V100 GPU, FlexRun achieves 2.69$\times$ higher performance on average for various GPT2 models. HLSTransform (he2024hlstransform, ) uses HLS to design a FPGA accelerator and synthesis combined with pipelining, memory unrolling, and memory partitioning and transfer optimizations, with the addition of 8-bit integer quantization (W8A8). On a tiny model with 110 million parameters, HLSTransform achieves 57.11 tokens/s on Xilinx Virtex UltraScale+ VU9P FPGA. SECDA-LLM (haris2024designing, ) utilizes quantization (W3A8) and designs an efficient FPGA-based LLM accelerators for the llama.cpp inference framework. By deploying on the PYNQ-Z1 board, it achieves 0.588 tokens/s for the TinyLlama model (1.1B). Chen et al. (chen2024understanding, ) investigate the feasibility and potential of model-speciic spatial acceleration for LLM inference on FPGAs. They introduce a comprehensive analytical model to estimate the LLM inference performance of FPGA accelerator with W4A8 quantization, and provide a library of high-level synthesis (HLS) kernels that are composable and reusable. For Llama2-7B, during prefilling phase, they can achieves about 213 tokens/s, 43 tokens/s, and 320 tokens/s on Xilinx Alveo U280, VCK5000, and VHK158 FPGAs, respectively. During decode stage, they can achieves about 200 tokens/s, 40 tokens/s, and 333 tokens/s on Xilinx Alveo U280, VCK5000, and VHK158 FPGAs, respectively.
3.1.5. ASIC
@@
Weight-Only Quantization. Despite the memory footprint reduction achieved by weight-only quantization, the actual computing performance is not really improved due to dequantization from integer to float. FIGNA (jang2024figna, ) proposes dedicated FP-INT arithmetic units designed specifically for FP-INT MAC operations and integrates them on the accelerator. FIGNA with FP16-INT4 provides 3.2768 TOPS computing power and 26.58W power consumption by considering all memory access in 28nm at 100MHz. Estimated result shows that for OPT-6.7B it can achieve 21.332 tokens/s in decode stage. Different from normal quantization methods, MECLA (qin2024mecla, ) proposes a parameter-efficient scaling sub-matrix partition method (SSMP) to decompose large weight matrices into several tiny-scale source sub-matrices (SS) and derived sub-matrices (DS). For memory issues, SSMP avoids accessing the full weight matrix but only requires small SS and DS scaling scalars. For computation issues, the proposed accelerator fully exploits the intermediate data reuse of matrix multiplication via on-chip matrix regrouping, inner-product multiplication re-association, and outer-product partial sum reuse. Totally, it can reduce 83.6% memory access and 72.2% computation. MECLA provides 14.008 TOPS computing power and $\sim$96W (1.9763W+94W) power consumption by considering all memory access under 28nm. For Llama2-7B and BLOOM-7B, compared to NVIDIA V100 GPU, MECLA achieves 6.74$\times$ and 5.91$\times$ inference speedup ($\sim$161 tokens/s and 141 tokens/s, respectively).
@@
Weight-Activation Quantization. Based on the key insight that outliers are important while the normal values next to them are not, OliVe (guo2023olive, ) adopts an outlier-victim pair (OVP) quantization and handles outlier values locally with low hardware overheads. This enables a memory-aligned W4A4/W8A8 quantization, which can be efficiently integrated to the existing hardware accelerators like systolic array and tensor core. OliVe provides 0.71 TOPS computing power and $\sim$8W (0.2806W+7.9872W) power consumption by considering all memory access under 22nm. Estimated results shows that for OPT-6.7B it can achieve 9.173 tokens/s in decode stage. Li et al. (li2024quantization, ) uniformly group weights and activations to ensure workload balance for hardware, and propose two approaches called channel sorting and channel selection to enhance the performance of quantization. It provides 1.43 TOPS computing power and $\sim$8.5W (0.472W+7.9872W) power consumption by considering all memory access under 65nm. Estimated results shows that for OPT-6.7B it can achieve 19.733 tokens/s in decode stage. Tender (10609625, ) decomposes weight and activation matrices by groups with different size to smooth the impact of outliers. And the format of scale factors are powers of two apart, which avoids explicit dequantization and extension to the commodity tensor compute hardware. It is 7.174W (1.60W+5.574W) power consumption by considering HBM2 memory access under 28nm. Result shows that for OPT-6.7B, Tender achieves 1.33$\times$ speedup (53.33 tokens/s) than NVIDIA A100 GPU.
3.1.6. PIM/NDP
@@
Figure 4. LLM ($\sim$ 7 billion parameters) decode stage throughput (batch size 1) vs power on different platforms with quantization.
3.1.7. Quantitative Comparison
We first compare the power consumption, inference speed and energy efficiency for different hardware platforms, in Figure 4. For quantization, power consumption ranges from 3W to 450W, with inference speeds between 3 tokens/s and 1998 tokens/s. The energy efficiency ranges from 0.0167 tokens/J to 46.66 tokens/J. T-MAC (wei2024t, ) (CPU) achieves the lowest power consumption with 3W and Guo et al. (guo2024towards, ) (PIM/NDP) achieves the highest throughput with batch size 1 (pre-silicon simulation result). And Guo et al. (guo2024towards, ) also achieves the highest energy efficiency with 46.66 tokens/J.
  • For CPUs, power consumption ranges from 3W to 385W, with inference speeds between 3 tokens/s and 50 tokens/s, located in the bottom part of the figure. The energy efficiency ranges from 0.0167 tokens/J to 2.38 token/J. Additionally, we observe edge CPUs (including CPU SoCs) with 3W to 6W power consumption exhibit higher energy efficiency (0.544 tokens/J to 2.38 tokens/J).
  • For GPUs, power consumption ranges from 40W to 450W, with inference speeds between 18 tokens/s and 194 tokens/s, situated in the upper right part of the figure. The energy efficiency ranges from 0.0667 tokens/J to 0.825 token/J. Compared to other hardware, GPUs can achieve higher absolute inference speeds due to their high computing power and high bandwidth. When quantization methods are used, the memory access bottlenecks in LLM inference are alleviated, further unlocking computing power.
  • For FPGAs, power consumption ranges from 45W to 225W, with inference speeds between 40 tokens/s and 333 tokens/s, also in the upper right part of the figure. The energy efficiency ranges from over 0.178 tokens/J to 1.85 tokens/J, which is higher than GPUs and server CPUs.
  • For ASICs, power consumption ranges from 6.3W to 96.66W, with inference speeds between 9.173 tokens/s and 161.086 tokens/s, found in the upper left section of the figure. The energy efficiency ranges from 0.803 tokens/J to over 7.434 tokens/J, outperforming CPU, GPU and FPGA hardware platforms.
  • For PIM/NDPs, power consumption ranges from 11.516W to 42.819W, with inference speeds between 481 tokens/s and 1998 tokens/s, found in the upper left section of the graph. The energy efficiency outperforms other hardware platforms.

### v4

@@ 3.1.3 GPU
Weight-Activation Quantization. In addition to hardware units that support FP16 computations, NVIDIA GPUs also provide hardware units that support INT4, INT8, and FP8 computations. The number of these computation units can be 2$\times$ and 4$\times$ greater than FP16 on each chip. Compared to weight-only quantization, weight-activation quantization can utilize INT4, INT8, and FP8 computations, thereby maximizing the peak computational performance of the GPU. Since the prefill phase in LLM inference is compute-bound, weight-activation quantization can significantly enhance performance during this stage. LLM.int8 [121] uses vector-wise quantization with separate normalization constants for each inner product in the matrix multiplication, to quantize most of the features. For the outliers, it isolates the outlier feature dimensions into a 16-bit matrix multiplication while still more than 99.9% of values are multiplied in 8-bit. For BLOOM-176B model, LLM.int8 achieves 4.05 tokens/s, 30.3 tokens/s and 109.77 tokens/s for batch size 1, 8 and 32, respectively, on 3 A100 GPUs in decode phase. The inference speed is slightly slower but close to 16-bit inference with less GPU consumption. SmoothQuant [122] enables 8-bit weight and 8-bit activation (W8A8) quantization for LLMs. Based on the fact that weights are easy to quantize while activations are not, SmoothQuant smooths the activation outliers by offline migrating the quantization difficulty from activations to weights with a mathematically equivalent transformation. SmoothQuant enables an INT8 quantization of both weights and activations for all the matrix multiplications in LLMs. SmoothQuant achieves up to 1.56$\times$ speedup and 2$\times$ memory reduction for LLMs with negligible loss in accuracy. QUIK [123] is for the first time, that the majority of inference computations for LLMs can be performed with both weights and activations being cast to 4 bits. QUIK compresses most of the weights and activations to 4-bit, while keeping some outlier weights and activations in higher-precision. It also provides GPU kernels matching the QUIK format with highly-efficient layer-wise runtimes, which lead to practical end-to-end throughput improvements of up to 3.4$\times$ relative to FP16 execution in prefill phase.
@@
Prevalent quantization schemes (e.g., W8A8) cannot fully leverage the capabilities of modern GPUs, such as 4-bit integer operators, resulting in sub-optimal performance. To maximize the throughput, Atom [124] significantly boosts serving throughput by using low-bit operators and considerably reduces memory consumption via low-bit quantization. It attains high accuracy by applying a novel mixed-precision and fine-grained quantization process. For single batch inference, Atom can achieve about 30 tokens/s for on a NVIDIA RTX 4090 GPU. Atom improves end-to-end throughput by up to 7.73$\times$ compared to the FP16 and by 2.53$\times$ compared to INT8 quantization, while maintaining the same latency target.
@@
Compared to integer quantization, floating-point (FP) quantization can better handle long-tail or bell-shaped distributions, and it has emerged as a default choice in many hardware platforms. LLM-FP4 [125] quantizes both weights and activations in LLMs down to 4-bit floating-point values (W4A4) with negligible accuracy loss. Due to the lack of PF4 computing unit in GPUs, its decoding speed maybe slower than FP16 baseline.
3.1.4 FPGA
@@
Weight-Activation Quantization. FlexRun [129] uses 8-bit quantization (W8A8), conducts an in-depth design space exploration to find the best accelerator architecture for a target LLM model, and automatically reconfigures the accelerator based on the exploration results. With the implementation on Intel Stratix 10 GX and MX FPGAs, FlexRun outperforms the current state-of-the-art FPGA-based accelerator by 1.15$\times$–1.50$\times$ for GPT2, respectively. Compared to Nvidia’s V100 GPU, FlexRun achieves 2.69$\times$ higher performance on average for various GPT2 models. HLSTransform [130] uses HLS to design a FPGA accelerator and synthesis combined with pipelining, memory unrolling, and memory partitioning and transfer optimizations, with the addition of 8-bit integer quantization (W8A8). On a tiny model with 110 million parameters, HLSTransform achieves 57.11 tokens/s on Xilinx Virtex UltraScale+ VU9P FPGA. SECDA-LLM [131] utilizes quantization (W3A8) and designs an efficient FPGA-based LLM accelerators for the llama.cpp inference framework. By deploying on the PYNQ-Z1 board, it achieves 0.588 tokens/s for the TinyLlama model (1.1B). Chen et al. [132] investigate the feasibility and potential of model-speciic spatial acceleration for LLM inference on FPGAs. They introduce a comprehensive analytical model to estimate the LLM inference performance of FPGA accelerator with W4A8 quantization, and provide a library of high-level synthesis (HLS) kernels that are composable and reusable. For Llama2-7B, during prefilling phase, they can achieves about 213 tokens/s, 43 tokens/s, and 320 tokens/s on Xilinx Alveo U280, VCK5000, and VHK158 FPGAs, respectively. During decode stage, they can achieves about 200 tokens/s, 40 tokens/s, and 333 tokens/s on Xilinx Alveo U280, VCK5000, and VHK158 FPGAs, respectively. On-Device Qwen2.5 [135] leverages AWQ with FPGA-accelerated execution pipelines and achieves 5.1 tokens/s for Qwen-0.5B on Xilinx Kria KV260 edge platform. TeLLMe [137] is the first ternary LLM accelerator for low-power FPGAs (e.g., AMD KV260) that fully supports both prefill and autoregressive decoding using 1.58-bit weights and 8-bit activations. Under a 7W power budget, it delivers up to 9 tokens/s throughput over 1024-token contexts and prefill latencies of 0.55-1.15s for 64-128 token prompts. LightMamba [136] proposes an W4A4 and W8A8 FPGA-friendly post-training quantization algorithm that features rotation-assisted quantization and power-of-two SSM quantization to reduce the majority of computation to 4-bit. It can achieve 93 tokens/s, 7.21 tokens/s and 3.61 tokens/s on Xilinx Alveo U280 FPGA and Versal VCK190 FPGA for Mamba2-2.7B. TerEffic [138] is aslo tailored for ternary-quantized LLM inference. Experimental results demonstrate that the HBM-assisted architecture processes 290 tokens/s and 727 tokens/s for Llama-7B and 2.7B parameter models while consuming only 46W on AMD Alveo U280. Li et al. [139] deploy a 4-bit LLaMA2-7B model, achieving a decoding speed of 4.9 token/s on Zynq-based KV260 with 6.57W. MEADOW [140] reduces the off-chip memory access for LLMs with a novel token-parallel head-sequential (TPHS) dataflow and mixed-precision quantization with weight packing. It achieves 0.25-2 tokens/s for OPT-1.1B on the low power Xilinx ZCU102 FPGA platform that consumes less than 10W. TTD [141] develops a tensor-train decomposition (TTD) for LLMs with a further hardware implementation on FPGA. The compressed LLMs are further implemented on FPGA hardware within a highly efficient group vector systolic array (GVSA) architecture, which has DSP-shared parallel vector PEs for TTD inference, as well as optimized data communication in the accelerator. Experimental results show that it achieves 69.7 tokens/s and 65.8 tokens/s for ChatGLM3-6B and Llama2-7B on AMD Alveo V80 FPGA, respectively. LlamaF [142] leverages W8A8 quantization techniques, asynchronous computation, and a fully-pipelined accelerator to enhance efficiency. The evaluations show that for TinyLlama-1.1B model on a Xilinx ZCU102 platform, it can achieve 1.5 tokens/s throughput with 0.291 tokens/J energy efficiency. LNS-LLM [143] points out that LLM exhibits significant outliers when directly adopting the Logarithmic Number System (LNS) format and presents a data-format/architecture co-design to bright this gap. On the format side, it proposes a dynamic and asymmetric LNS format, and on the architecture side, realizes the dynamic LNS format in a systolic array. Experimental results show that on an Alveo U280 FPGA as a prototype, LNS-LLM can effectively handle the outliers and resolve the mismatch between LNS and LLM.
3.1.5 ASIC
@@
Weight-Only Quantization. Despite the memory footprint reduction achieved by weight-only quantization, the actual computing performance is not really improved due to dequantization from integer to float. FIGNA [146] proposes dedicated FP-INT arithmetic units designed specifically for FP-INT MAC operations and integrates them on the accelerator. FIGNA with FP16-INT4 provides 3.2768 TOPS computing power and 26.58W power consumption by considering all memory access in 28nm at 100MHz. Estimated result shows that for OPT-6.7B it can achieve 21.332 tokens/s in decode stage. Different from normal quantization methods, MECLA [147] proposes a parameter-efficient scaling sub-matrix partition method (SSMP) to decompose large weight matrices into several tiny-scale source sub-matrices (SS) and derived sub-matrices (DS). For memory issues, SSMP avoids accessing the full weight matrix but only requires small SS and DS scaling scalars. For computation issues, the proposed accelerator fully exploits the intermediate data reuse of matrix multiplication via on-chip matrix regrouping, inner-product multiplication re-association, and outer-product partial sum reuse. Totally, it can reduce 83.6% memory access and 72.2% computation. For Llama2-7B and BLOOM-7B, compared to NVIDIA V100 GPU, MECLA achieves 6.74$\times$ and 5.91$\times$ inference speedup ($\sim$161 tokens/s and 141 tokens/s, respectively) under 91.84W (32 processors) power consumption.
FineQ [152] quantizes 3 weights into 8-bit with 2/3-bit mixed-precision and introduces an accelerator utilizing temporal coding that effectively supports the quantization algorithm while simplifying the multipliers in the systolic array. It can achieve up to 1.79× energy efficiency and reduces the area of the systolic array by 61.2%. Anda [157] proposes an adaptive data format with group-shared exponent bits and dynamic mantissa bit allocation.
@@
Weight-Activation Quantization. Based on the key insight that outliers are important while the normal values next to them are not, OliVe [148] adopts an outlier-victim pair (OVP) quantization and handles outlier values locally with low hardware overheads. This enables a memory-aligned W4A4/W8A8 quantization, which can be efficiently integrated to the existing hardware accelerators like systolic array and tensor core. OliVe provides 0.71 TOPS computing power and 5.95W power consumption by considering all memory access under 22nm. Estimated results shows that for Llama-7B it can achieve 9.95 tokens/s in decode stage. Li et al. [149] uniformly group weights and activations to ensure workload balance for hardware, and propose two approaches called channel sorting and channel selection to enhance the performance of quantization. It provides 1.43 TOPS computing power and 2.36W power consumption by considering all memory access under 65nm. Estimated results shows that for Llama-7B it can achieve 20.04 tokens/s in decode stage. Tender [150] decomposes weight and activation matrices by groups with different size to smooth the impact of outliers. And the format of scale factors are powers of two apart, which avoids explicit dequantization and extension to the commodity tensor compute hardware. It is 7.174W (1.60W+5.574W) power consumption by considering HBM2 memory access under 28nm. Result shows that for OPT-6.7B, Tender achieves 1.33$\times$ speedup (53.33 tokens/s) than NVIDIA A100 GPU.
Cai et al. [151] propose a grouped adaptive two-range quantization (ATRQ) with an in-group embedded identifier to encode outliers and develop a low-overhead ATRQ decoder and an outlier-bitsplit PE to reduce the hardware overhead associated with high-bit-width outliers. Experimental results show that their design can achieve the similar speed (9.95 tokens/s) with OliVe for OPT-6.7B with $\sim$1.632W. OFQ-LLM [154] is an outlier-flexing quantization accelrator for LLM inference. It can achieves the similar speed with OliVe [148] with 1.64W for 7B models. M-ANT [155] proposes an efficient real-time quantization mechanism and implements a specific processing element (PE) to efficiently support MANT and incorporate a real-time quantization unit. By integrating these components into a systolic array, MANT unifies the group-wise weight and KV cache quantization and addresses the associated challenges. Evaluation shows it can achieve, on average, 1.96$\times$ speedup and 1.54$\times$ energy reduction to OliVe accelerator (19.502 tokens/s, 7.57W). Unlike previous bit-serial designs, which also provide flexibility but at the cost of performance due to its bit-wise temporal processing nature, FlexiBit’s architecture enables bit-parallel processing of any precision and format without compute unit under utilization [156]. AirGun [158] applies hardware efficient per-tensor quantization for non-sensitive modules and per-input channel quantization for sensitive modules, and proposes an accelerator that fully utilizes the advantages of AirGun (12.935 tokens/s, 6.1964W). Ecco [159] combines group-wise and nonuniform quantization with pre-defined shared k-means patterns and Huffman coding to exploit the inherent entropy characteristics of LLM cache data. It also introduces a parallel Huffman-based decoding process with a multi-stage pipeline design, reducing latency by two orders of magnitude and achieving throughput comparable to GPU L2 caches. OwL-P [160] leverages outliers and shared exponents to facilitate the compression of both model weights and activations. FGMP [161] proposes a fine grained FP4/FP8 mixed-precision quantization and a hardware implementation for performing dot products between two FP4 blocks, two FP8 blocks, or one FP4 and one FP8 block. LightRot [162] proposes a lightweight rotation scheme and dedicated hardware accelerator designed for 4-bit LLM inference.
3.1.6 PIM/NDP
@@
Through on-the-fly vector encoding and horizontal matrix layout, MVDRAM [167] eliminates the overheads introduced by the Processing using DRAM’s (PUD) fundamental limitation of column-to-column data movement. It can achieve 2-10 tokens/s and 7-20 tokens/J for Llama2-7B with 2-8-bit quantization.
ROMA [169] is a QLoRA accelerator with a hybrid storage architecture that uses ROM for quantized base models and SRAM for LoRA weights and KV cache. ROMA can achieve 11600-24100 tokens/s for 2-bit Llama3-8B under 33.1W power consumption.
Figure 4: LLM ($\sim$ 7 billion parameters) decode stage throughput (batch size 1) vs power on different platforms with quantization.
3.1.7 Quantitative Comparison
We first compare the power consumption, inference speed and energy efficiency for different hardware platforms, in Figure 4. For quantization, power consumption ranges from 3W to 450W, with inference speeds between 3 tokens/s and 2700 tokens/s. The energy efficiency ranges from 0.0057 tokens/J to 46.66 tokens/J. T-MAC [102] (CPU) achieves the lowest power consumption with 3W and WaferLLM [252] (ASIC) achieves the highest throughput with batch size 1 (pre-silicon simulation result). And Guo et al. [163] also achieves the highest energy efficiency with 46.66 tokens/J.
  • For CPUs, power consumption ranges from 3W to 385W, with inference speeds between 3 tokens/s and 50 tokens/s, located in the bottom part of the figure. The energy efficiency ranges from 0.0167 tokens/J to 2.38 token/J. Additionally, we observe edge CPUs (including CPU SoCs) with 3W to 9W power consumption exhibit higher energy efficiency (1 tokens/J to 2.38 tokens/J).
  • For GPUs, power consumption ranges from 40W to 450W, with inference speeds between 18 tokens/s and 343.4 tokens/s, situated in the upper right part of the figure. The energy efficiency ranges from 0.067 tokens/J to 0.859 token/J. Compared to other hardware, GPUs can achieve higher absolute inference speeds due to their high computing power and high bandwidth. When quantization methods are used, the memory access bottlenecks in LLM inference are alleviated, further unlocking computing power.
  • For FPGAs, power consumption ranges from 2.490W to 225W, with inference speeds between 3.61 tokens/s and 333 tokens/s, also in the upper right part of the figure. The energy efficiency ranges from over 0.154 tokens/J to 6.30 tokens/J, which is higher than GPUs and server CPUs.
  • For ASICs, power consumption ranges from 1.632W to 91.84W, with inference speeds between 9.95 tokens/s and 161.1 tokens/s, found in the upper left section of the figure. The energy efficiency ranges from 1.16 tokens/J to over 8.49 tokens/J, outperforming CPU, GPU and FPGA hardware platforms.
  • For PIM/NDPs, power consumption ranges from 11.516W to 70W, with inference speeds between 10 tokens/s and 1998 tokens/s, found in the upper left section of the graph. The energy efficiency outperforms other hardware platforms.